On July 1, 2019, Crane Company purchased new equipment for $75,000. Its estimated useful life was 5 years with a $8,000 salvage value. On December 31, 2022, the company estimated that the equipment’s remaining useful life was 10 years, with a revised salvage value of $5,000.

Compute the revised annual depreciation on December 31, 2022.

Revised annual depreciation $
Answer #1

Accumulated depreciation (July 1,2019 to Dec 31,2021) = (75000-8000/5*2.5) = 33500

Revised annual Depreciation = (75000-33500-5000)/10 = 3650 per year